OpenStudy (anonymous):

A crowbar is used as a lever. The effort force of 40 newtons moves 3 meters. The resistance force of 54 newtons moves 2 meters. What is the efficiency of the lever? 100% because energy must be conserved 90% 25% 18%

OpenStudy (anonymous):

well the work done pushing down on the lever is W_d = F*d = 40N * 3m =120Nm=120J the work done on the resistance is equal to: W_r = F*d = 54 N* 2 m = 108 Nm = 108 J You put in 120 J of work and only get out 108 therefore your crowbar is not 100% efficient efficiency = 100% * work_out / work_in = 100% * 108J / 120J = 90%